# TopSeg: A Multi-Scale Topological Framework for Data-Efficient Heart Sound Segmentation

### v1

@@ Abstract
Deep learning approaches for heart-sound (PCG) segmentation built on time–frequency features can be accurate but often rely on large expert-labeled datasets, limiting robustness and deployment. We present TopSeg, a topological representation-centric framework that encodes PCG dynamics with multi-scale topological features and decodes them using a lightweight temporal convolutional network (TCN) with an order- and duration-constrained inference step. To evaluate data efficiency and generalization, we train exclusively on PhysioNet 2016 dataset with subject-level subsampling and perform external validation on CirCor dataset. Under matched-capacity decoders, the topological features consistently outperform spectrogram and envelope inputs, with the largest margins at low data budgets; as a full system, TopSeg surpasses representative end-to-end baselines trained on their native inputs under the same budgets while remaining competitive at full data. Ablations at 10% training confirm that all scales contribute and that combining $H_{0}$ and $H_{1}$ yields more reliable S1/S2 localization and boundary stability. These results indicate that topology-aware representations provide a strong inductive bias for data-efficient, cross-dataset PCG segmentation, supporting practical use when labeled data are limited.
@@ 1 Introduction
Heart disease is the leading cause of mortality worldwide, responsible for approximately 17.9 million deaths annually [1], with auscultation serving as a primary, cost-effective diagnostic tool in clinical practice [2]. Accurate segmentation of phonocardiogram (PCG) signals into their primary components—first (S1) and second (S2) heart sounds, and the intervening systolic and diastolic intervals—aids physicians in cardiac diagnosis by helping identify timing abnormalities, detect murmurs, and assess valvular function [3]. Fig. 1 illustrates the segmentation task, where PCG signals are segmented into S1, S2, systolic, and diastolic phases.

### v2

@@ Abstract
Deep learning approaches for heart-sound segmentation built on time–frequency features can be accurate but often rely on large expert-labeled datasets, limiting robustness and deployment. We present TopSeg, a topological representation-centric framework that encodes Phonocardiogram (PCG) dynamics with multi-scale topological features and decodes them using a lightweight temporal convolutional network (TCN) with an order- and duration-constrained inference step. To evaluate data efficiency and generalization, we train exclusively on PhysioNet 2016 dataset with subject-level subsampling and perform external validation on CirCor dataset. Under matched-capacity decoders, the topological features consistently outperform spectrogram and envelope inputs, with the largest margins at low data budgets; as a full system, TopSeg surpasses representative end-to-end baselines trained on their native inputs under the same budgets while remaining competitive at full data. Ablations at 10% training confirm that all scales contribute and that combining $H_{0}$ and $H_{1}$ yields more reliable S1/S2 localization and boundary stability. These results indicate that topology-aware representations provide a strong inductive bias for data-efficient, cross-dataset PCG segmentation, supporting practical use when labeled data are limited.
@@ 1 Introduction
Heart disease is the leading cause of mortality worldwide, responsible for approximately 17.9 million deaths annually [21], with auscultation serving as a primary, cost-effective diagnostic tool in clinical practice [14]. Accurate segmentation of phonocardiogram (PCG) signals into their primary components—first (S1) and second (S2) heart sounds, and the intervening systolic and diastolic intervals—aids physicians in cardiac diagnosis by helping identify timing abnormalities, detect murmurs, and assess valvular function [18, 23]. Fig. 1 illustrates the segmentation task, where PCG signals are segmented into S1, S2, systolic, and diastolic phases.
@@ 5 Conclusion
6 Acknowledgment
This work was funded by Basic Research Program of Jiangsu Province under Grant BG2024027.